a mechanical oscilator has initial energy Enot= 50j having a damping coefficient b=1s. Calculate the time required to decrease to E= Enot/e


1
Expert's answer
2022-03-27T18:41:21-0400

Answer

Energy is given

"E=E_0e^{-bt\/m}"

Energy is given at time t is "E=\frac{E_0}{e}"

"E_0/e=E_0/e^{t/m}
\\ \frac{1}{e}=\frac{1}{e^(t/1)}"

Now calculate

"e=e^{t\/1}"

So t=1sec.